# Non-linear degenerate parabolic flow equations and a finer differential structure on Wasserstein spaces

### v1

@@ 3.7.2 Grönwall type estimates and Hausdorff–continuity of the set of solutions
For $(M,g)=(\mathbb{R}^{n},g^{eucl})$, the result follows by immediate induction using (3.58) and (3.59). For an arbitrary manifold with bounded geometry, Nash’s embedding theorem allows to conclude. $\square$
@@
Let $p>2$ and $T>0$. Let $\mathcal{D}\in\mathcal{W}_{p}(M)$ be a dense subset. With the asumptions of Theorem 3.44, suppose that for all $\mu_{0}\in\mathcal{D}$ and all $\frac{1}{p}$–Hölder–continuous $\mu:[0,T]\to\mathcal{W}_{p}(M)$ with $\mu(0)=\mu_{0}$, the linear MDE (3.125) has a unique smooth solution (See Definition 3.28) in $\mathcal{W}_{p}(M)$. Then, for all $\nu_{0}\in\mathcal{W}_{p}(M)$, there exists a unique solution $\nu:[0,T]\to\mathcal{W}_{p}(M)$ of (3.14) with initial value $\nu_{0}$ which is a uniform limit of Average Flow Approximation Series (See Definition 3.27) in $\mathcal{W}_{p}(M)$ on $[0,T]$.
@@
By assumption, for all $\mu_{0}\in\mathcal{D}$, $S(V,T,\mu_{0})$ is a singleton. Hence, by Theorem 3.44, $S(V,T,\nu_{0})$ is a singleton for all $\nu_{0}\in\mathcal{W}_{p}(M)$. $\square$
@@
Lemma 3.47.
Assume $(M,g)$ is of bounded geometry. Let $p\geq 2$, $T>0$, $R>0$ and $\mu_{0}\in\mathcal{W}_{p}(M)$. Let $\bigl{(}\mathcal{}\mathfrak{P}(\frac{1}{N})\bigr{)}_{N\in\mathbb{N}}$ be a family of partitions of $[0,T]$. There exists a constant $\mathcal{K}>0$, such that for any two continuous curves $V,W:[0,T]\mapsto\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$ uniformly supported in the ball with radius $R$, there exists $M\in\mathbb{N}$ depending on $\mathcal{W}_{2}^{\infty}(V,W)$, such that for all $N\geq M$ and all $t\in[0,T]$, the respective LAFAS denoted by $\mu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}$ and $\nu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}$ satisfy
$\mathcal{W}_{2}(\mu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(t),\nu^{\mathcal{}\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(t))_{\infty}\leq\mathcal{K}\sqrt{t}\mathcal{W}_{2}^{\infty}(V,W).$ (3.136)
Let $\mathcal{S}(V,T,\mu_{0})$ be the set of smooth solutions of (3.125) in $\mathcal{W}_{p}(M)$, with initial value $\mu_{0}\in\mathcal{W}_{p}(M)$ on $[0,T]$. Then, the map
$\mathcal{S}:\mathcal{C}^{0}\bigl{(}[0,T],\mathcal{W}_{2}(\mathcal{X}^{2,\infty}(M))\bigr{)}$ $\to\mathbb{P}\biggl{(}\mathcal{C}^{0,\frac{1}{p}}\bigl{(}[0,T],\mathcal{W}_{2}(M)\bigr{)}\biggr{)}$ (3.137)
$V$ $\mapsto S(V,T,\mu_{0}),$
is Lipschitz with respect to the Hausdorff–distance on $\mathbb{P}\biggl{(}\mathcal{C}^{0,\frac{1}{p}}\bigl{(}[0,T],\mathcal{W}_{2}(M)\bigr{)}\biggr{)}$.

### v2

@@ 3.7.2 Grönwall type estimates and Hausdorff–continuity of the set of solutions
For $(M,g)=(\mathbb{R}^{n},g^{eucl})$, the result follows by immediate induction using (3.57) and (3.58). For an arbitrary manifold with bounded geometry, we use induction in Lemma 3.34. $\square$
Lemma 3.47.
Assume $(M,g)$ is of bounded geometry. Let $p\geq 2$, $T>0$, $R>0$ and $\mu_{0}\in\mathcal{W}_{p}(M)$. Denote by
  • $S_{p}(V,T,\mu_{0})$ the set of smooth solutions of (LMDE) in $\mathcal{W}_{p}(M)$ for a curve $V:[0,T]\to\mathcal{W}_{p+2}(\mathcal{X}^{2,\infty}(M))$, with initial value $\mu_{0}\in\mathcal{W}_{p}(M)$ on $[0,T]$,
  • $\mathcal{B}_{p}\coloneqq 2^{\frac{p+3}{2}}3\sqrt{p-1}$.
Let $V,W:[0,T]\mapsto\mathcal{W}_{p+2}(\mathcal{X}^{2,\infty}(M))$ be two curves of vector–field probabilities satisfying $\mathcal{M}_{p+2}(V_{t}),\mathcal{M}_{p+2}(W_{t})\leq R^{p+2}$ for all $t\in[0,T]$. Let $\bigl(\mathcal{}\mathfrak{P}(\frac{1}{N})\bigr)_{N\in\mathbb{N}}$ be a family of partitions of $[0,T]$. Denote by $\bigl(\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}\bigr)_{N\in\mathbb{N}}$ and $\bigl(\nu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}\bigr)_{N\in\mathbb{N}}$ the respective (LAFAS) with initial value $\mu_{0}$ associated to $V$ and $W$ respectively. Then, there exists a constant $C(M)>0$ depending only on $M$, with $C(\mathbb{R}^{n})=1$ for all $n\in\mathbb{N}$, such that the following hold:
  • Suppose $p=2$. There exists $M\in\mathbb{N}$, such that for all $N\geq M$ and for all $t\in\bigl[0,\frac{1}{2(R+1)^{2}}\bigr]$,
    $\mathcal{W}_{2}\bigl(\mu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(t),\nu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(t)\bigr)\leq C(M)6\mathcal{W}_{2}(V,W)_{\infty}\sqrt{t}.$ (3.131)
  • Suppose $p>2$. There exist $\mathfrak{S}_{p}>0$, and for all $2<q\leq 4$, an integer $M_{q}\in\mathbb{N}$, such that for all $N\geq M_{q}$ and for all $t\in\bigl[0,\frac{\mathfrak{S}_{p}}{(R+1)^{2}}\bigr]$,
    $\mathcal{W}_{p}\bigl(\mu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(t),\nu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(t)\bigr)\leq C(M)\mathcal{B}_{p}\mathcal{W}_{p}(V,W)_{\infty}t^{\frac{1}{q}}.$ (3.132)
  • In particular, for all $p>2$, $\mu\in S_{p}(V,T,\mu_{0})$, $\nu\in S_{p}(W,T,\mu_{0})$, and $t\in\bigl[0,\frac{\mathfrak{S}_{p}}{(R+1)^{2}}\bigr]$,
    $\mathcal{W}_{p}\bigl(\mu_{t},\nu_{t}\bigr)\leq C(M)\mathcal{B}_{p}\mathcal{W}_{p}(V,W)_{\infty}\sqrt{t}.$ (3.133)
Let $p>2$. Furthermore, there exists a constant $\mathcal{K}>0$, depending only on $R$, $T$, $p$ and $M$, such that for all $2<q\leq 4$, there exists $M_{q}\in\mathbb{N}$, such that for all $N\geq M_{q}$ and all $t\in[0,T]$,
$\mathcal{W}_{p}(\mu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(t),\nu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(t))\leq\mathcal{K}\mathcal{W}_{p}(V,W)_{\infty}t^{\frac{1}{q}},$ (3.134)
and
$\mathcal{W}_{p}(\mu_{t},\nu_{t})\leq\mathcal{K}\mathcal{W}_{p}(V,W)_{\infty}\sqrt{t}.$ (3.135)
In particular, the map
$\mathcal{S}_{p}:\mathcal{C}^{0}\bigl([0,T],\mathcal{W}_{p+2}(\mathcal{X}^{2,\infty}(M))\bigr)$ $\to\mathbb{P}\biggl(\mathcal{C}^{0,\frac{1}{p}}\bigl([0,T],\mathcal{W}_{p}(M)\bigr)\biggr)$ (3.136)
$V$ $\mapsto S_{p}(V,T,\mu_{0}),$
is locally Lipschitz–continuous with respect to the Hausdorff–distance on the power set $\mathbb{P}\biggl(\mathcal{C}^{0,\frac{1}{p}}\bigl([0,T],\mathcal{W}_{p}(M)\bigr)\biggr)$.
Proof.
Let us first work with $(M,g)=(\mathbb{R}^{n},g^{eucl})$ and $p=2$. Let $\varepsilon>0$. Denote for all $N\in\mathbb{N}$, $(x_{N,i})_{i\in I_{N}}\coloneqq\mathcal{}\mathfrak{P}(\frac{1}{N})$ in increasing order.
Proof of the first point: First, we prove by induction that there exists $\overline{M}\in\mathbb{N}$, such that for all $t\in\bigl[0,\frac{1}{2(R+1)^{2}}\bigr]$ and all $N\geq\overline{M}$,
$\mathcal{W}_{2}\bigl(\mu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(t),\nu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(t)\bigr)^{2}\leq 36\mathcal{W}_{2}(V,W)_{\infty}^{2}t.$ (3.137)
Initialisation: First, (3.59) and (3.60) in Lemma 3.33 yield that there exists $M\in\mathbb{N}$—depending on $\mathcal{W}_{2}(V,W)_{\infty}$—, such that for all $N\geq M$,
[…]
Hypothesis: Now, let $i\in I_{N}$ such that $x_{N,i}\leq\frac{1}{2(R+1)^{2}}$ and suppose that
$\mathcal{W}_{2}\bigl(\mu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(x_{N,i}),\nu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(x_{N,i})\bigr)^{2}\leq 36x_{N,i}\mathcal{W}_{2}(V,W)^{2}_{\infty}.$
Step: Then, by (3.59), there exists $\overline{M}\geq M$, such that for all $N\geq\overline{M}$,
$\mathcal{W}_{2}\bigl(f^{V(x_{N,i})}_{x_{N,i+1}-x_{N,i}}(\mu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(x_{N,i})),f^{W(x_{N,i})}_{x_{N,i+1}-x_{N,i}}(\nu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(x_{N,i}))\bigr)^{2}$
$\qquad\leq\mathcal{W}_{2}\bigl(\mu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(x_{N,i}),\nu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(x_{N,i})\bigr)^{2}$
$\hskip 18.49988pt+2(1+\varepsilon)(x_{N,i+1}-x_{N,i})\biggl(9\mathcal{W}_{2}(V,W)^{2}_{\infty}$ (3.139)
$\hskip 18.49988pt+3R\mathcal{W}_{2}\bigl(\mu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(x_{N,i}),\nu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(x_{N,i})\bigr)\mathcal{W}_{2}(V,W)_{\infty}\biggr)$
$\qquad\leq 36x_{N,i}\mathcal{W}_{2}(V,W)^{2}_{\infty}$
$\hskip 18.49988pt+2(1+\varepsilon)\mathcal{W}_{2}(V,W)^{2}_{\infty}\bigl(9+18(R+1)\sqrt{x_{N,i}}\bigr)(x_{N,i+1}-x_{N,i})$
$\qquad\leq 36x_{N,i+1}\mathcal{W}_{2}(V,W)^{2}_{\infty},$
since $x_{N,i}\leq\frac{1}{2(R+1)^{2}}$. Thus, we conclude by induction.
Proof of the second point: For $p>2$, the statement follows similarly from equations (3.55) and (3.56). Let $2<q\leq 4$.
Initialisation: There exists $M_{q}\in\mathbb{N}$—depending on $\mathcal{W}_{p}(V,W)_{\infty}$—such that there exists $\mathcal{K}>0$ depending only on $p$, such that for all $N\geq M$,
$\mathcal{W}_{p}\bigl(\mu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(x_{N,1}),\nu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(x_{N,1})\bigr)^{p}\leq\mathcal{K}\mathcal{W}_{p}(V,W)^{2}_{\infty}x_{N,1}^{\frac{p}{2}}$ (3.140)
$\hskip 18.49988pt\leq\mathcal{B}_{p}^{p}\mathcal{W}_{p}(V,W)_{\infty}^{p}x_{N,1}^{\frac{p}{q}}.$
Hypothesis: Now, let $i\in I_{N}$ and suppose that
$\mathcal{W}_{p}\bigl(\mu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(x_{N,i}),\nu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(x_{N,i})\bigr)^{p}\leq\mathcal{B}_{p}^{p}x_{N,i}^{\frac{p}{q}}\mathcal{W}_{p}(V,W)^{p}_{\infty}.$
Step: One readily shows that the induction hypothesis implies that there exits $\overline{M_{q}}\geq M_{q}$, such that for all $N\geq\overline{M_{q}}$,
$\mathcal{W}_{p}\bigl(\mu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(x_{N,i+1}),\nu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(x_{N,i+1})\bigr)^{p}\leq\mathcal{W}_{p}(V,W)_{\infty}^{p}\biggl(\mathcal{B}_{p}^{q}x_{N,i}$ (3.141)
$\hskip 18.49988pt+2^{p}q\biggl(9(p-1)\mathcal{B}_{p}^{q-2}x_{N,i}^{\frac{q-2}{q}}+pR^{2}\mathcal{B}_{p}^{q}x_{N,i}$
$\hskip 18.49988pt\qquad+6R\mathcal{B}_{p}^{q-1}x_{N,i}^{\frac{q-1}{q}}\biggr)(x_{N,i+1}-x_{N,i})\biggr)^{\frac{p}{q}}$
$\hskip 18.49988pt\leq\mathcal{W}_{p}(V,W)_{\infty}^{p}\mathcal{B}_{p}^{p}\biggl(x_{N,i}$
$\hskip 18.49988pt+\biggl(\frac{1}{2}x_{N,i}^{\frac{q-2}{q}}+2^{p+2}\bigl(pR^{2}x_{N,i}+6Rx_{N,i}^{\frac{q-1}{q}}\bigr)\biggr)(x_{N,i+1}-x_{N,i})\biggr)^{\frac{p}{q}}.$
In particular, there exists a constant $\mathfrak{S}_{p}>0$ depending only on $p$, such that if $x_{N,i}\leq\frac{\mathfrak{S}_{p}}{(R+1)^{2}}$,
$\mathcal{W}_{p}\bigl(\mu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(x_{N,i+1}),\nu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(x_{N,i+1})\bigr)^{p}\leq\mathcal{B}_{p}^{p}\mathcal{W}_{p}(V,W)_{\infty}^{p}x_{N,i+1}^{\frac{p}{q}}.$ (3.142)
Proof of the global estimates for (LAFAS): Now, another proof by induction using equations (3.55) and (3.56) and a subsequent Grönwall–type argument imply the existence of $\mathcal{K}>0$, depending only on $R$, $T$, $p$ and $M$, such that for all $2<q\leq 4$, there exists $M_{q}\in\mathbb{N}$, such that for all $N\geq M_{q}$ and all $t\in[0,T]$,
$\mathcal{W}_{p}(\mu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(t),\nu^{\mathcal{}\mathfrak{P}\bigl(\frac{1}{N}\bigr)}(t))_{\infty}\leq\mathcal{K}t^{\frac{1}{q}}\mathcal{W}_{p}(V,W)_{\infty}.$ (3.143)
Proof of the estimates for smooth solutions: Let $\mu\in S_{p}(V,T,\mu_{0})$. By Theorem 3.45, there exists a sequence of partitions $\bigl(\mathcal{}\mathfrak{P}(\frac{1}{N})\bigr)_{N\in\mathbb{N}}$ such that $t\mapsto\mu(t)$ is a uniform limit of $\bigl(\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}\bigr)_{N\in\mathbb{N}}$. Since we assume $p>2$, Theorem 3.45 yields that there exists an increasing sequence $N_{k}$ such that $\bigl(\nu^{\mathcal{}\mathfrak{P}(\frac{1}{N_{k}})}\bigr)_{N\in\mathbb{N}}$ converges uniformly to a $\nu\in S_{p}(W,T,\mu_{0})$. The uniform convergence implies that all the above estimates also hold or $\mu$ and $\nu$. In particular, for all $2<q\leq 4$ and all $t\in\bigl[0,\frac{\mathfrak{S}_{p}}{(R+1)^{2}}\bigr]$,
$\mathcal{W}_{p}(\mu_{t},\nu_{t})\leq\mathcal{B}_{p}t^{\frac{1}{q}}\mathcal{W}_{p}(V,W)_{\infty}.$ (3.144)
Therefore,
$\mathcal{W}_{p}(\mu_{t},\nu_{t})\leq\mathcal{B}_{p}t^{\frac{1}{2}}\mathcal{W}_{p}(V,W)_{\infty}.$ (3.145)
In a similar fashion, we show that there exists $\mathcal{K}>0$, depending only on $R$, $T$, $p$ and $M$, such that for all $t\in[0,T]$,
$\mathcal{W}_{p}(\mu_{t},\nu_{t})\leq\mathcal{K}t^{\frac{1}{2}}\mathcal{W}_{p}(V,W)_{\infty}.$ (3.146)
Thus, for every smooth solution of (LMDE) for $V$, there exists a smooth solution of (LMDE) for $W$, such that
$\mathcal{W}_{p}(\mu,\nu)_{\infty}\leq\mathcal{K}T^{\frac{1}{2}}\mathcal{W}_{p}(V,W)_{\infty},$ (3.147)
and vice–versa. This yields the local Lipschitz regularity of $\mathcal{S}_{p}$.
The results for an arbitrary manifold with bounded geometry follow analogously by Lemma 3.34. $\square$
3.7.3 Reduction of the non–linear setting to the linear setting
Introducing the linear equation (LMDE) allows us to derive the uniqueness of solutions of (MDEd) which are limits of AFAS from the uniqueness of solutions of (LMDE) which are limits of LAFAS using a fixed point argument. This argument is detailed in our second main result, Theorem 3.48.
Theorem 3.48.
Let $p>2$, $T>0$ and $\mu_{0}\in\mathcal{W}_{p}(M)$. Assume:
  • $(M,g)$ is of bounded geometry.
  • $V$ is a probability vector–field which satisfies:
    • $V$ is Lipschitz with respect to the pp–Wasserstein distance.
    • There exists a constant $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]\in\mathcal{W}_{p+2}(\mathcal{X}^{2,\infty}(M))$ and $\mathcal{M}_{p+2}(V[\mu])\leq R^{p+2}$.
  • For all $\frac{1}{p}$–Hölder–continuous $\mu:[0,T]\to\mathcal{W}_{p}(M)$ with $\mu(0)=\mu_{0}$, the linear MDE (LMDE) has a unique smooth solution (See Definition 3.44) for $t\mapsto V[\mu]$ with initial value $\mu_{0}$ in $\mathcal{W}_{p}(M)$.
Then:
[…]
  • For any sequence of partitions of the interval $[0,T]$, $\biggl(\mathcal{}\mathfrak{P}(\frac{1}{N})\biggr)_{N\in\mathbb{N}}$ which satisfies $\lim_{N\to\infty}\mathfrak{p}(\mathcal{}\mathfrak{P}(\frac{1}{N}))=0$, the sequence of Average Flow Approximation Series $\biggl(\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}\biggr)_{N\in\mathbb{N}}$ with initial value $\mu_{0}$ converges uniformly to this unique smooth solution in $\mathcal{W}_{p}(M)$.
Proof.
Let $T>0$ and $\bigl(\mathcal{}\mathfrak{P}(\frac{1}{N})\bigr)_{N\in\mathbb{N}}$ be a sequence of partitions of $[0,T]$. Let $\mathcal{S}:\mathcal{C}^{\frac{1}{p}}([0,T],\mathcal{W}_{p}(M))\to\mathcal{C}^{\frac{1}{p}}([0,T],\mathcal{W}_{p}(M))$ be the map sending a $\frac{1}{p}$–Hölder–continuous path $\mu:[0,T]\to\mathcal{W}_{p}(M)$ to the unique smooth solution of the corresponding linear MDE (LMDE) with initial value $\mu_{0}$. We will show that $\mathcal{S}$ is a contraction with respect to the $\mathcal{W}_{p}$–distance for $T$ small enough.
Let $c_{1},c_{2}\in\mathcal{C}^{\frac{1}{p}}([0,T],\mathcal{W}_{p}(M))$. Denote by $t\mapsto\mu(t)$ and $t\mapsto\nu(t)$ the smooth solutions of (LMDE) for $V\circ c_{1}$ and $V\circ c_{2}$ respectively. By assumption, the linear MDEs (LMDE) for $V\circ c_{1}$ and $V\circ c_{2}$ have unique smooth solutions. Thus, for any sequence of partitions $\bigl(\mathcal{}\mathfrak{P}(\frac{1}{N})\bigr)_{N\in\mathbb{N}}$, the Linear Average Flow Approximation Series $\bigl(\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}\bigr)_{N\in\mathbb{N}}$ and $\bigl(\nu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}\bigr)_{N\in\mathbb{N}}$ converge uniformly to those solutions with respect to the pp–Wasserstein distance by Theorem 3.45.
Hence, Lemma 3.47 yields that there exists a constant $\mathcal{K}>0$, depending only on $p$, $T$, $R$ and $M$, such that for all $t\in[0,T]$,
$\mathcal{W}_{p}(\mu(t),\nu(t))\leq\mathcal{K}t^{\frac{1}{2}}\mathcal{W}_{p}(V\circ c_{1},V\circ c_{2})_{\infty}.$ (3.148)
In particular,
$\mathcal{W}_{p}(\mu,\nu)_{\infty}\leq L\mathcal{K}T^{\frac{1}{2}}\mathcal{W}_{p}(c_{1},c_{2})_{\infty}.$ (3.149)
Therefore, choosing $T$ small enough, ensures that $\mathcal{S}$ is a contraction. Thus, Banach’s fixed point theorem yields that $\mathcal{S}$ has a unique fixed point $s:[0,T]\to\mathcal{W}_{p}(M)$. Hence, (MDEd) has a unique solution $s:[0,T]\to\mathcal{W}_{p}(M)$ with initial value $\mu_{0}$ on $[0,T]$, which is a limit of Linear Average Flow Approximation Series associated to $s$.
Let us conclude by showing that $s$ is also the unique solution of (MDEd) in $\mathcal{W}_{p}(M)$ with initial value $\mu_{0}$, which is a limit of Average Flow Approximation Series. Suppose there exist two such solutions $s_{1}$ and $s_{2}$. Hence, there exist $\bigl(\mathcal{}\mathfrak{P}_{1}(\frac{1}{N})\bigr)_{N\in\mathbb{N}}$ and $\bigl(\mathcal{}\mathfrak{P}_{2}(\frac{1}{N})\bigr)_{N\in\mathbb{N}}$, two sequences of partitions of $[0,T]$ and Average Flow Approximation Series $s^{\mathcal{}\mathfrak{P}_{1}(\frac{1}{N})}_{1}$ and $s^{\mathcal{}\mathfrak{P}_{2}(\frac{1}{N})}_{2}$ converging uniformly in $\mathcal{W}_{p}(M)$ respectively to $s_{1}$ and $s_{2}$ with respective steps given by $\mathcal{}\mathfrak{P}_{1}(\frac{1}{N})$ and $\mathcal{}\mathfrak{P}_{2}(\frac{1}{N})$. Let $\varepsilon>0$. Since the convergence is uniform and $V$ is LL–Lipschitz with respect to the pp–Wasserstein distance, there exists $N_{0}\in\mathbb{N}$, such that for all $n\geq N_{0}$,
Now, it suffices to show that this implies that the LAFAS $\mathfrak{s}^{\mathcal{}\mathfrak{P}_{1}(\frac{1}{N})}_{1}$ and $\mathfrak{s}^{\mathcal{}\mathfrak{P}_{2}(\frac{1}{N})}_{2}$ associated to the linear MDEs (LMDE) for $s_{1}$ and $s_{2}$ also converge to $s_{1}$ and $s_{2}$ respectively. This follows again by induction using (3.150) and (3.55) in Lemma 3.33. Thus, $s_{1}$ and $s_{2}$ are both fixed points of $\mathcal{S}$ and, hence, $s_{1}=s_{2}$.
@@ 3.7.3 Reduction of the non–linear setting to the linear setting
Let $p>2$ and $T>0$. Let $\mathcal{D}\in\mathcal{W}_{p}(M)$ be a dense subset. With the assumptions of Theorem 3.46, suppose that for all $\mu_{0}\in\mathcal{D}$ and all $\frac{1}{p}$–Hölder–continuous $\mu:[0,T]\to\mathcal{W}_{p}(M)$ with $\mu(0)=\mu_{0}$, the linear MDE (LMDE) has a unique smooth solution (see Definition 3.30) in $\mathcal{W}_{p}(M)$. Then, for all $\nu_{0}\in\mathcal{W}_{p}(M)$, there exists a unique solution $\nu:[0,T]\to\mathcal{W}_{p}(M)$ of (MDEd) with initial value $\nu_{0}$ which is a uniform limit of Average Flow Approximation Series (see Definition 3.29) in $\mathcal{W}_{p}(M)$ on $[0,T]$.
@@
By Theorem 3.48, for all $\mu_{0}\in\mathcal{D}$, $S(V,T,\mu_{0})$ is a singleton. Hence, by Theorem 3.46, $S(V,T,\nu_{0})$ is a singleton for all $\nu_{0}\in\mathcal{W}_{p}(M)$. $\square$
@@
3.7.4 Strong Generalised Central Limit Theorem and uniqueness of smooth solutions
Now, we are ready to prove our third and last main result, the uniqueness of smooth solutions of (MDEd). First, we prove that the smooth solutions of (LMDE) are unique for sufficiently regular initial values. Then, we combine Theorem 3.48 and some of our Grönwall type estimates to conclude.
Proposition 3.51.
Let $R\geq 1$, $p>2$, $\mu_{0}\in\mathcal{C}_{c}^{\infty}(M)$, and $V:[0,T]\mapsto\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$ be a $\frac{1}{p}$–Hölder–continuous curve such that for all $t\in[0,T]$, $V_{t}\in\mathcal{W}_{p+2}(\mathcal{X}^{2,\infty}(M))$ and $\mathcal{M}_{p+2}(V_{t})\leq R^{p+2}$. Then there exists a unique smooth solution of equation (LMDE) for $V$ with initial value $\mu_{0}$ in $\mathcal{W}_{p}(M)$.